Fermi National Laboratory


Accelerator Update

Wednesday February 25
The day shift began with a Tevatron (TeV) expert conducting end-of-store studies with the remains of store 3254, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:10 AM, Operations terminated store 3254.

Operators prepared the Collider Detector at Fermilab (CDF) and D-zero (D0) collision halls for access at 8:23 AM. Experimenters were out of both halls by 9:50 AM.

Cryo system technicians completed their repairs on the TeV sector B1 wet engine at 9:40 AM.

Operations put the TeV through a dry squeeze at 11 AM and then into shot setup at 11:25 AM.

At 2:35 PM, Operations established store 3256 with an initial luminosity of 59.68E30. Operations stashed some protons into the Recycler. Pbar resumed stacking at 2:44 PM.

Operations reported that with the help of a Switchyard (SY) expert, the Meson Test (MTest) fixed target experiment began taking beam at 3:38 PM.

Switchyard experts completed their studies at 7 PM and MTest experimenters finished using their beam at 8:23 PM.

Thursday February 26
The midnight shift began with Operations monitoring store 3256, with MiniBooNE taking beam, and with Pbar stacking.

At 3:41 AM, the Main Injector (MI) RF suffered from a vacuum burst. Everything recovered okay.

At 6:36 AM, a Booster RF station (BRF4) tripped several times. Two other stations, BRF1 and BRF8, also tripped. Operators got these stations running, but called in a tech to repair BRF4.

A SY expert began a study period at 8:24 AM. Operations established beam to MTest at 8:46 AM.

At 9:50 AM, a TeV expert began (now don't laugh) a helix enlargement study. The study was finished by 10:15 AM, but the enlarged helix was left in the machine.

Operations turned off Booster beam at 10 AM so techs could conduct RF work. Power supply techs performed maintenance work during this time. All work was completed within an hour.

Pbar resumed stacking at 11:02 AM.

At 11:05 AM, Booster experts began a collimator study that kept beam intensity very low. Operations turned off MiniBooNe during this study.

A Linac quadrupole power supply (L:QPS229) tripped off at 11:22 AM. An expert had the power supply back up by 11:35 AM.

At 12:06 PM, the TeV suffered a six-house quench due to a vacuum problem at sector A4.

Operations prepared the CDF and D0 collision halls for access at 1:12 PM.

Safety system techs began testing the Pbar safety system at 1:43 PM.

Vacuum techs said they had to access TeV sector A4. Operators began preparing the TeV for access at 1:53 PM.

Safety techs began testing the MI and MiniBooNE systems at 2 PM.

Pbar resumed stacking at 3:13 PM and then began a reverse proton study at 3:33 PM.

SY began a ramp study and MI began a 2.5MHz study at 3:53 PM.

At 7:08 PM, vacuum techs completed their TeV work. The A4 cryo system recovery began.

The Pbar reverse proton study finished at 8:10 PM.

Cryo techs reported having problems cooling down the TeV A4 cryo system due to poor vacuum. Operations called in vacuum techs to investigate.

Pbar resumed stacking at 10 PM.

SY experts completed their studies and Operations halted all beam to the Meson area at 10:53 PM.

Friday February 27
The midnight shift began with Pbar stacking, with Recycler and MiniBooNE experts conducting studies, and with the TeV in access to investigate a vacuum leak.

At 12:46 AM, the vacuum techs came out of the TeV. They installed more vacuum pumps.

At 1:02 AM, the cryo techs began cool down of the A4 cryo system. The cool down was completed by 3 AM. Operations began the TeV startup.

Operations reported that the TeV A4 beam-valves closed at 4:38 AM. The vacuum looked okay.

Operators prepared the TeV to access and investigate a possible problem with a vacuum gauge device at 5:30 AM. Techs were out of the TeV by 6:42 AM.

Pbar resumed stacking at 7:06 AM.

At 7:09 AM, a Booster RF station (BRF12) tripped off on a ground fault indication. Operators contacted an expert.

Operations again began the TeV turn-on at 7:22 AM.

Today's Plans
The plans for today are to complete the TeV turn-on and then establish a new store. The Pbar stack is presently at 212E10 particles.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/27/2004   email Fermilab